# A Multi-Fidelity Parametric Framework for Reduced-Order Modeling using Optimal Transport-based Interpolation: Applications to Diffused-Interface Two-Phase Flows

### 04 Mar 2026 v1

@@ 6.1. Rider-Kothe vortex
A classical test case for interface capturing techniques consists in the deformation of a two-dimensional circular bubble by a shear flow [58]. In particular, in the classical test, a droplet of radius $R=0.15$ centered at $(0.0,0.25)$ in a $0.5,0.5]^{2}$ square periodic domain is advected by the following divergence-free velocity field:
$\begin{bmatrix}u_{1}\\ u_{2}\end{bmatrix}=\begin{bmatrix}\sin^{2}(\pi(x_{1}+0.5))\sin(2\pi(x_{2}+0.5))\cos\left(\frac{\pi t}{T}\right)\\ -\sin(2\pi(x_{1}+0.5))\sin^{2}(\pi(x_{2}+0.5))\cos\left(\frac{\pi t}{T}\right)\end{bmatrix}$, (33)
where $T$ denotes the characteristic period of the shear flow, which we have chose as the classical value $T=4$ for this particular test. Under the action of such a velocity field, in the first half of the period, the initial circular bubble is strongly deformed into a thin filament. After velocity reversal, the filament is stretched back to the initial condition. It is already clear that this benchmark possesses several challenges from both the high-fidelity and reduced-order modeling perspectives. Indeed, the strong deformation of the interface leading to sharp and very localized features, the advection-dominated nature of the problem with significant variations of its support, and the complexity of the nonlinear dynamics, all contribute to making this physical phenomena a very challenging test case to verify the capabilities of our strategies in accurately predicting the phase field.
Thus, given the highly refined grid needed to capture the sharp and thin features of the interface, and the notable differences between snapshots at different time instants, standard ROM techniques may not be a suitable choice, while displacement interpolation is expected to provide consistent and accurate predictions, generating synthetic but physically meaningful approximations at a negligible computational cost. For this problem, we consider a $3^{\mathrm{rd}}$-order polynomial discretization on a series of increasingly finer meshes from $128^{2}$ to $512^{2}$ degrees of freedom. Time integration was performed using a classical $4^{\mathrm{th}}$ order Runge-Kutta scheme.
Moreover, to test the generalizability of the proposed procedure to the parameterized setting, we consider the parametric extension of the test case by changing the initial radius of the bubble in the following way:
@@
An overview of the problem’s dynamics corresponding to different values for the initial radius $R$ and two different level of refinement (low-fidelity with $256^{2}$ and high-fidelity with $512^{2}$) is shown in Figure 5. If not specified otherwise, we will always consider the case $R=0.11$ to test our framework for the Rider-Kothe vortex.
(a) $R=0.1$.
(b) $R=0.15$.
(c) $R=0.2$.
(d) $R=0.1$.
(e) $R=0.15$.
(f) $R=0.2$.
Figure 5. Low-fidelity (top) and high-fidelity (bottom) phase-field isocontours for different initial conditions of the Rider-Kothe vortex at $t=T/2$.
@@ 6.1.1. Proper Orthogonal Decomposition and Displacement Interpolation
Before presenting the performance of the novel strategies proposed in this work, we want to further motivate and numerically show the main reasons to exploit displacement interpolation instead of a classical POD-based reduced-order models. Here, we aim to highlight the limitations of standard linear reduction strategies in reconstructing the temporal evolution of the phase field in two-phase flows. Towards this goal, we consider the aforementioned Rider-Kothe vortex problem and compare the results obtained with POD projection of the high-fidelity solutions and via displacement interpolation.
@@
We remark that we have introduced the nomenclature checkpoints for the set of snapshots used to perform the pairwise displacement interpolation. Thus, to provide a fair comparison with POD, the number of checkpoints used in the displacement interpolation will always match the number of snapshots available for the POD projection. Indeed, the checkpoints are the only data nominally available by the OT procedure as well as by the POD.
@@
Figure 6 shows an example of the predicted phase field obtained exploiting the two strategies (POD and DI) with $N_{c}=30$ checkpoints. It is evident that performing displacement interpolation between high-fidelity data produces more accurate results than using the POD basis. This expected improvement is indeed related to the spatial differences in the support regions between snapshots: the POD projection essentially computes a linear combination of the available modes leading to a noticeable smearing of the interface and intermediate unphysical values for the reconstructed field. In contrast, displacement interpolation correctly predicts values of either zero or one in the whole spatial domain, detecting with high accuracy the evolution of the moving feature and its sharp variation.
@@
Figure 6. Phase field at $t\approx 0.2$ for the Rider-Kother vortex ($256\times 256$). Left, FOM; center, DI ($N_{c}=30$); right, POD projection ($N_{c}=30$).
@@
In Figures 7 and 8, instead, we present a more quantitative evaluation of the relative errors of displacement interpolation and POD projection varying the total number of checkpoints. We notice that, for a small number of checkpoints, the relative errors show only minor differences between the two approaches. However, as the number of checkpoints increases, displacement interpolation between consecutive data points provides significantly more accurate results compared to POD projection. In particular, note the clearly intermittent behavior at initial and late stages of the dynamics. In these regions, where the droplet moves at its highest speed, the POD projection struggles to accurately capture the dynamics of the problem, and provides seemingly better results only in reconstructing the states that have been used to construct the POD basis. Although displacement interpolation yields only modest improvements compared to POD projection in terms of relative error (see Figure 8), we emphasize that this observation is largely influenced by the specific error metric adopted. Such a metric may be partially misleading, as it does not fully capture the key features of interest in the solution, namely the interface morphology and the accuracy of the reconstructed phase-field values.
@@
Figure 7. Relative error for DI (left) and POD projection (right) for the Rider-Kother vortex ($256\times 256$).
@@
Figure 8. Mean relative error for DI and POD projection for the Rider-Kother vortex ($256\times 256$).
For these reasons, in Figure 9, we analyze the problem from a more modeling-oriented perspective by examining the total area where $0.1<\phi_{1}<0.9$, that in the case of bounded values, provides an approximation of the droplet’s surface.
@@
Specifically, Figure 9 shows that the general trend of the FOM behaves as expected for this problem: the droplet is initially circular, then stretches into a thin ligament, maximizing the total surface area, and finally returns toward its initial shape. The displacement interpolation approach captures this behavior accurately, while, in contrast, POD projection deviates significantly from the expected trend, particularly at the initial and final stages of the dynamics, where the prescribed advection velocity reaches its maximum.
(a) DI
(b) POD
@@
Figure 9. Normalised area of the spatial domain such that $0.1<\phi_{1}<0.9$ using DI (left) and POD projection (right).
@@
Furthermore, even in regions where common support exists, the POD projection essentially acts as a high-order interpolation, and even increasing the number of basis spanning the POD space, the benefits in accuracy are counteracted by the introduction of spurious oscillations near sharp features. This behavior is clearly visible in Figure 10, where we show a one-to-one comparison of the upper bounds of the phase field predicted by displacement interpolation and POD projection. It is evident that DI consistently remains very close to the FOM simulations, whereas the POD projection exhibits strong oscillations linked to unphysical behavior around the correct values.
(a) DI
(b) POD
@@
Figure 10. Maximum value of the predicted phase field as a function of time using DI (left) and POD projection (right).
@@ 6.1.2. MF-OT-ROM
In the previous section, we highlighted how POD-based ROMs perform poorly for advection-dominated problems, such as those arising in numerical simulations of two-phase flows. While POD can achieve reasonable results in terms of relative error, more case-specific metrics, such as boundedness and interfacial area, reveal significant deviations from the correct dynamics.
In this section, we consider the proposed extensions of the OT-based surrogate modeling for the simulation of two-phase flows. We begin with the MF-OT-ROM approach introduced previously, where displacement interpolation is used to interpolate the residuals between high- and low-fidelity solutions. The interpolated residual is then employed to correct the low-fidelity simulations in order to recover, as accurately as possible, an approximation of the high-fidelity solutions.
For this test case, we consider three levels of fidelity: $128\times 128$, $256\times 256$, and $512\times 512$, computing the pairwise residuals between the different grid resolutions.
As in the previous section, to perform a complete study on the performance of the strategy with respect to the available data, we vary the number of checkpoints used for the displacement interpolation. Figures 11 and 12 quantify the relative errors produced by the proposed approach, measured as the difference between the low-fidelity simulation (corrected or not via OT) and the high-fidelity simulation. In Figure 11, we show the relative error between low- and high-fidelity solutions as a function of time, varying the number of checkpoints. The uncorrected low-fidelity solution is shown in black, while a color gradient from blue to red indicates an increasing number of checkpoints. With relatively few checkpoints (e.g., $N_{c}=10$), the corrected low-fidelity solution can locally increase the error relative to the high-fidelity solution, although the mean error decreases (see Figure 12). As the number of checkpoints increases, the relative error decreases throughout the whole trajectory, demonstrating progressively better agreement with the high-fidelity solution.
Figure 11. Relative error between low- and high-fidelity solutions w.r.t. different number of checkpoints for Rider-Kothe vortex with $R=0.11$. Horizontal dashed lines represent the mean error for each temporal trajectory. Low-fidelity, $256\times 256$; high-fidelity $512\times 512$.
In Figure 12, we consider both settings with $128\times 128$ and $256\times 256$ exploited as low-fidelity, and we investigate the decay of the mean error. Notice that the trend is almost identical between the two cases, but always over-performing the uncorrected approximations, suggesting that the correction is effective for both resolutions and that in both cases we are accurately capturing all the features from the dynamics. Indeed, by reporting the relative errors, the improvement obtained by correcting the $128\times 128$ solution to match the $256\times 256$ solution is, in relative terms, almost the same as that obtained by correcting the $256\times 256$ solution to match the $512\times 512$ solution.
@@
From a qualitative perspective, Figures 13 and 14 show the isolines $\phi_{1}=0.5$, representing the interface location, predicted by the corrected low-fidelity solutions, along with the low-fidelity (dashed black) and high-fidelity (dash-dotted black) references. In Figure 13, the $128\times 128$ solution is considered as low-fidelity, whereas in Figure 14, the $256\times 256$ solution serves as the low-fidelity reference, and increasing the number of checkpoints significantly improves in both cases the predicted interface location.
Figure 13. Isocontour of $\phi=0.5$ for $256\times 256$ HF (dash-dotted black), $128\times 128$ LF (dashed black) and corrected LFs (color gradient) at $t=1.95$.
Figure 14. Isocontour of $\phi=0.5$ for $512\times 512$ HF (dash-dotted black), $256\times 256$ LF (dashed black) and corrected LFs (color gradient) at $t=1.95$.
In Figure 15, we show the full predicted field for $\phi_{1}$ near the top of the droplet at $t=1.95$. The first row displays the high-fidelity simulation, while the second row shows the low-fidelity resolution, where, from left to right, the amount of OT correction is increased: the left panel shows the uncorrected low-fidelity solution, while the center and right panels show the corrected low-fidelity solutions using $N_{c}=10$ and $N_{c}=80$ checkpoints, respectively. The nice performance of the methodology is already evident exploiting only $10$ checkpoints, with the solution showing a significant improvement in terms of interface thickness. Finally, using $N_{c}=80$ produces a corrected solution that is nearly indistinguishable from the high-fidelity simulation.
Figure 15. Comparison between $512\times 512$ HF simulations (repeated on the top row) and different levels of $256\times 256$ LF in the bottom row at $t=1.95$. Specifically, from left to right: Uncorrected LF, Corrected LF with $N_{c}=10$ and Corrected LF with $N_{c}=80$.
@@ 6.1.3. PMF-OT-ROM
In this section, we increase the complexity of the discussed benchmark by considering a fully parametric setting. In this scenario, two different strategies can be used to correct a low-fidelity simulation for a given parameter when no high-fidelity solution is readily available. The first approach consists of computing the residual from neighboring trajectories, interpolating this residual in parameter space, and then performing a second interpolation in time, reverting to the MF-OT-ROM approach presented earlier. The second strategy, instead, involves applying displacement interpolation directly to high-fidelity data from neighboring trajectories to generate a new high-fidelity trajectory for unseen parameters, which can then be used to correct the corresponding low-fidelity solution. We will focus on this latter approach for the remainder of the work as it has shown better stability properties in the resolution of the OT problem.
In Figure 16 we report the relative error between the corrected low-fidelity solution and the validation high-fidelity solution for an unseen parametric trajectory. We can observe that increasing the number of checkpoints significantly decreases the errors throughout the whole dynamics. In Figure 17 we show the mean error over the entire trajectory, which further highlights the decrease in error as the number of checkpoints increases.
Figure 16. Relative error between low- and high-fidelity for different number of checkpoints for the Rider-Kothe vortex with unseen parameter $R=0.11$. Horizontal dashed lines represent the mean error for each temporal trajectory. Low-fidelity, $128\times 128$; high-fidelity $256\times 256$.
Figure 17. Mean relative errors against number of checkpoints for the unseen parameter $R=0.11$. Low-fidelity, $128\times 128$; high-fidelity $256\times 256$.
Finally, in Figure 18, as a metrics which is more specific for this problem, we investigate the upper bound and total interface’s surface by comparing high-fidelity, uncorrected low-fidelity and a series of corrected low fidelities with the proposed approach with different number of checkpoints. Notice that as the number of checkpoints increase we obtain an almost perfect match in both quantities.
(a) Upper bound.
(b) Interfacial area.
Figure 18. Maximum value of the predicted phase field as a function of time (left) and interfacial area (right). Dash-dotted line, HF; dashed, uncorrected LF; color gradient, corrected LFs. Low-fidelity, $128\times 128$; high-fidelity $256\times 256$.

### 27 Mar 2026 v2

@@ 7.1. Rider-Kothe vortex
A classical test case for interface capturing techniques consists in the deformation of a two-dimensional liquid column by a shear flow [72]. In particular, in the classical test, a droplet of radius $R=0.15$ centered at $(0.5,0.75)$ in a $0,1]^{2}$ square periodic domain is advected by the following divergence-free velocity field:
$\textbf{v}=\begin{bmatrix}v_{1}\\[2.84544pt] v_{2}\end{bmatrix}=\begin{bmatrix}\sin^{2}(\pi x_{1})\sin(2\pi x_{2})\cos\left(\frac{\pi t}{T}\right)\\[2.84544pt] -\sin(2\pi x_{1})\sin^{2}(\pi x_{2})\cos\left(\frac{\pi t}{T}\right)\end{bmatrix}$, (32)
where $T$ denotes the characteristic period of the shear flow, which we have chosen as the classical value $T=4$ for this particular test. Under the action of such a velocity field, during the first half of the period the initial circular bubble is strongly deformed into a thin filament. After the velocity reversal, the filament is advected back toward its initial configuration.
This benchmark already highlights several challenges from the full-order perspective. At the full-order level, the strong deformation of the interface represents a well-known difficulty for interface-capturing techniques, which should ideally preserve the interface as sharp as possible while satisfying boundedness requirements and maintaining numerical stability.
Challenges also emerge from a model order reduction standpoint, as the presence of a sharp interface moving rapidly across the computational domain is a typical feature of advection-dominated flows. As a result, the need for highly refined grids to resolve thin interface structures, combined with the pronounced variability between snapshots at different time instances, can significantly affect the performance of conventional ROM approaches. By contrast, displacement interpolation is anticipated to deliver consistent and accurate predictions, producing synthetic yet physically meaningful approximations at a negligible computational cost.
For this problem, we consider as the Full Order Model (FOM) a $3^{\mathrm{rd}}$-order polynomial spatial discretization on a series of increasingly finer meshes from $128^{2}$ to $512^{2}$ degrees of freedom, while we performed time integration using a classical $4^{\mathrm{th}}$ order Runge-Kutta scheme. For this first test case, we consider a dataset consisting of $401$ snapshots, corresponding to a time window of $\Delta\tau=0.01$. The time step sizes used in the simulations are approximately $\Delta t_{128}\approx 10^{-3}$, $\Delta t_{256}\approx 5\cdot 10^{-4}$, and $\Delta t_{512}\approx 2.5\cdot 10^{-4}$ for the three spatial resolutions, to satisfy CFL constraints.
Moreover, we propose the following parametric extension of the test case by changing the initial radius of the bubble as:
@@
An overview of the problem’s dynamics corresponding to different values for the initial radius $R$ and two different levels of refinement (LF with $256^{2}$ and HF with $512^{2}$) is shown in Figure 5. If not specified otherwise, we will always consider the case $R=0.11$ to test our framework for the Rider-Kothe vortex in the non-parametric scenario.
(a) $R=0.1$
(b) $R=0.15$
(c) $R=0.2$
(d) $R=0.1$
(e) $R=0.15$
(f) $R=0.2$
Figure 5. LF (top) and HF (bottom) phase-field isocontours for different initial conditions of the Rider-Kothe vortex at $t^{*}=T/2$.
@@ 7.1.1. Proper Orthogonal Decomposition and Displacement Interpolation
Before presenting the performance of the novel strategies proposed in this work, we want to further motivate and numerically show the main reasons to exploit displacement interpolation instead of a classical POD-based reduced-order models for advection-dominated flows. Here, we aim to highlight the limitations of standard linear reduction strategies in reconstructing the temporal evolution of the phase-field in two-phase flows. Towards this goal, we consider the aforementioned Rider-Kothe vortex problem and compare the results obtained with POD projection and displacement interpolation of the HF solutions.
@@
We remark that we have introduced the nomenclature checkpoints for the set of snapshots used to perform the pairwise displacement interpolation. Thus, to provide a fair comparison with POD, the number of checkpoints used in the displacement interpolation will always match the number of snapshots available for the POD projection. Indeed, the checkpoints are the only data nominally available to both OT- and POD-based methodologies.
@@
Figure 6 shows an example of the predicted phase-field obtained by exploiting the two strategies (POD and DI) with $N_{c}=30$ checkpoints. It is evident that performing displacement interpolation between HF data produces more accurate results than using the POD basis. This expected improvement is indeed related to the spatial differences in the support regions between snapshots: the POD projection essentially computes a linear combination of the available modes leading to a noticeable smearing of the interface, large oscillations in the union of the supports, and intermediate unphysical values for the reconstructed field. In contrast, displacement interpolation correctly predicts sharply varying values between zero and one in the whole spatial domain, detecting with high accuracy the evolution of the interface.
@@
Figure 6. Phase-field at $t\approx 0.2$ for the Rider-Kother vortex at resolution $256^{2}$ with FOM (left), DI (center) and POD (right) for $N_{c}=30$.
@@
In Figures 7 and 8, we present a more quantitative evaluation of the relative errors of displacement interpolation and POD projection varying the total number of checkpoints. In particular, for this analysis we select the number of checkpoints so that they are uniformly spaced across the available snapshots, enabling a fair comparison between DI, POD, and the FOM solutions. Since the dataset contains $401$ snapshots in total, this condition is satisfied for $N_{c}=11,21,41,51,81$. We notice that for a small number of checkpoints, the relative errors show only minor differences between the two approaches. However, as the number of checkpoints increases, displacement interpolation between consecutive data points provides significantly more accurate results compared to POD projection. In particular, note the clearly intermittent behavior at the initial and late stages of the dynamics (see Figure 7). In these regions, where the droplet moves at its highest speed, the POD projection struggles to accurately capture the dynamics of the problem, and provides seemingly better results only in reconstructing the states that have been used to construct the POD basis. From Figure 8, depicting the analysis of the decay of the error w.r.t. the number of checkpoints $N_{c}$, it may seem that displacement interpolation yields only modest improvements compared to POD projection in terms of relative error, but we emphasize that this observation is largely influenced by the specific error metric adopted, which is biased towards pointwise discrepancy methodologies. Indeed, such a metric is partially misleading, as it does not fully capture the key features of interest in the solution, namely the interface morphology and the accuracy of the reconstructed phase-field values.
@@
Figure 7. Relative error behavior w.r.t. time for the Rider-Kother vortex at resolution $256^{2}$ with DI (left) and POD projection (right).
@@
Figure 8. Mean relative error w.r.t. the number of checkpoints $N_{c}$ for the Rider-Kother vortex at resolution $256^{2}$ with DI and POD projection.
@@
For these reasons, we also analyze the problem from a more modeling-oriented perspective by examining the total area where $0.1<\phi<0.9$ that, in the case of bounded values, provides an approximation of the droplet’s interfacial area. Specifically, Figure 9 shows that the general trend of the FOM behaves as expected for this problem: the droplet is initially circular, then stretches into a thin ligament, maximizing the total surface area, and finally comes back to its initial shape due to the periodicity of the advection field. Displacement interpolation captures this behavior accurately, whereas, in contrast, POD projection deviates significantly from the expected trend, particularly at the initial and final stages of the dynamics, where the prescribed advection velocity reaches its maximum magnitude.
@@
Figure 9. Normalized area of the spatial domain such that $0.1<\phi<0.9$ using DI (left) and POD projection (right) for different $N_{c}$ values.
@@
Furthermore, even in regions where common support exists, the POD projection essentially acts as a high-order interpolation, and even increasing the number of basis spanning the POD space, the benefits in accuracy are counteracted by the introduction of spurious oscillations near sharp features. This behavior is clearly visible in Figure 10, where we show a one-to-one comparison of the upper bounds of the phase-field predicted by displacement interpolation and POD projection. It is evident that DI consistently remains very close to the FOM simulations, whereas the POD projection exhibits strong oscillations around the unitary upper bound.
@@
Figure 10. Maximum value of the predicted phase-field as a function of time using DI (left) and POD projection (right) for different $N_{c}$ values.
@@ 7.1.2. MF-OT-ROM
Having highlighted how POD-based ROMs perform poorly when applied to two-phase flows, especially when evaluated on more case-specific metrics such as boundedness and interfacial area, we are ready to present the performance of the proposed extensions of OT-based surrogate modeling. We begin with the MF-OT-ROM approach introduced previously, where displacement interpolation is used to interpolate the residuals between high- and low-fidelity solutions. The interpolated residual is then employed to correct the LF solutions in order to recover, as accurately as possible, an approximation of the HF solutions. For this test case, we consider three levels of fidelity: $128^{2}$, $256^{2}$ and $512^{2}$, computing the pairwise residuals between the different grid resolutions.
As in the previous section, to perform a complete accuracy study of the strategy with respect to the available data, we vary the number of checkpoints used for displacement interpolation. Figures 11 and 12 quantify the relative errors produced by the proposed approach, measured as the difference between the LF solution (corrected or not via OT) and the HF solution. In Figure 11, we show the relative error between low- and high-fidelity solutions ($256^{2}$ and $512^{2}$) as a function of time, varying the number of checkpoints. The uncorrected LF solution is shown in black, while a color gradient from blue to red indicates an increasing number of checkpoints. With relatively few checkpoints (e.g., $N_{c}=10$), the corrected LF solution can locally increase the error relative to the HF solution, although the mean error decreases (see Figure 12). As the number of checkpoints increases, the relative error decreases throughout the whole trajectory, demonstrating progressively better agreement with the HF solution.
Figure 11. Relative error between low- and high-fidelity solutions ($256^{2}$ and $512^{2}$) w.r.t. different number of checkpoints for Rider-Kothe vortex with $R=0.11$. Horizontal dashed lines represent the mean error for each temporal trajectory.
In Figure 12, we consider both settings with $128^{2}$ and $256^{2}$ exploited as LF, and we investigate the decay of the mean error. Notice that the trend is almost identical between the two cases, but always over-performing the uncorrected approximations, suggesting that the correction is effective for both resolutions and that in both cases we are accurately capturing all the features from the dynamics. Indeed, by reporting the relative errors, the improvement obtained by correcting the $128^{2}$ solution to match the $256^{2}$ solution is, in relative terms, almost the same as that obtained by correcting the $256^{2}$ solution to match the $512^{2}$ solution.
@@
From a qualitative perspective, Figures 13 and 14 show the isolines $\phi=0.5$, representing the interface location, predicted by the corrected LF solutions, along with the uncorrected LF (dashed black) and HF (dash-dotted black) references. In Figure 13, the $128^{2}$ solution is considered as LF, whereas in Figure 14, the $256^{2}$ solution serves as the LF reference, and increasing the number of checkpoints significantly improves in both cases the predicted interface location, correctly recovering tail and bubble structures. This analysis confirms the importance of exploiting refined meshes for the simulation of two-phase flows, and thus further motivates the need for multi-fidelity corrections to obtain efficient simulations.
Figure 13. Isocontour of $\phi=0.5$ for $256^{2}$ HF (dash-dotted black), $128^{2}$ LF (dashed black) and corrected LFs (color gradient) at $t=1.95$.
Figure 14. Isocontour of $\phi=0.5$ for $512^{2}$ HF (dash-dotted black), $256^{2}$ LF (dashed black) and corrected LFs (color gradient) at $t=1.95$.
In Figure 15, we show the predicted field for $\phi$ near the top of the droplet at $t=1.95$. From left to right, the amount of OT correction is increased: the left panel shows the uncorrected LF solution, the two center panels show the corrected LF solutions using $N_{c}=10$ and $N_{c}=80$ checkpoints, respectively, while the right panel show the HF solution. The nice performance of the methodology is already evident exploiting only $10$ checkpoints, with the solution showing a significant improvement in terms of interface thickness, while using $N_{c}=80$ produces a corrected solution that is nearly indistinguishable from the HF simulation.
Figure 15. Phase-field isocontours at $t=1.95$. From left to right: uncorrected LF ($256^{2}$), corrected LF with $N_{c}=10$, corrected LF with $N_{c}=80$, and reference HF ($512^{2}$).
@@ 7.1.3. PMF-OT-ROM
In this section, we increase the complexity of the Rider-Kothe vortex by considering its parametric extension. Indeed, HF solutions may not be directly available for all parametric trajectories, making the problem more challenging but also more relevant for applications where new parametric dynamics must be computed efficiently. In this context, we propose to apply displacement interpolation to HF data from neighboring trajectories in order to generate a new synthetic HF trajectory for previously unseen parameter values (see the discussion in Section 5 concerning the computational cost). Such synthetic checkpoints can then be used to compute the residual and correct the corresponding LF solution.
In Figure 16 we report the relative error between the corrected LF solution and the validation HF solution for an unseen parametric trajectory. We can observe that increasing the number of checkpoints significantly decreases the errors throughout the whole dynamics. In Figure 17 we show the mean error over the entire trajectory, which further highlights the decrease in error as the number of checkpoints increases.
Figure 16. Relative error between low- and high-fidelity ($128^{2}$ and $256^{2}$) w.r.t. different number of checkpoints for the Rider-Kothe vortex with unseen parameter $R=0.11$. Horizontal dashed lines represent the mean error for each temporal trajectory.
Figure 17. Mean relative errors against number of checkpoints for the unseen parameter $R=0.11$ with LF $128^{2}$ and HF $256^{2}$.
Finally, in Figure 18, as a metric which is more specific for this problem, we investigate the upper bound and total interfacial area by comparing HF, uncorrected LF and a series of corrected LFs with the proposed approach with different number of checkpoints. Notice that as the number of checkpoints increases we obtain an almost perfect match in both quantities.
(a) Upper bound
(b) Interfacial area
Figure 18. Maximum value of the predicted phase-field as a function of time (left) and interfacial area (right). Dash-dotted line, HF $256^{2}$; dashed, uncorrected LF $128^{2}$; color gradient, corrected LFs.